What is the effect on the interference pattern observed in a Young’s double slit experiment when the width of the slits is doubled? Give reasons for your answer.

Grade:12

1 Answers

Pramod J AskiitiansExpert-IIT-B
13 years ago

Dear student,

 

 

 

\begin{figure} \begin{center} \leavevmode \epsfxsize=2.8 in \epsfbox{fig22-8.eps}\end{center}\end{figure}

 

$\displaystyle\Delta$y = $\displaystyle{\frac{\lambda L}{d}}$

 

when the slit distance is increased distance between fringes increases